Can you please help me with this word problem? I have been working on this problem forever and yet to get an answer.
Several hours after departure the two ships described to the right are 340 miles apart. If the ship traveling south traveled 140 miles farther than the other, how many miles did they each travel?
How far did the eastbound ship travel? ________ miles
How far did the southbound ship travel? ________ miles

Answer by stanbon(75887)

Several hours after departure the two ships described to the right are 340 miles apart. If the ship traveling south traveled 140 miles farther than the other, how many miles did they each travel?
------------------
You have a right triangle with legs x and "x+140" and hypotenuse 340 miles
----
Solve:
x^2 + (x+140)^2 = 340^2
---
x^2 + x^2 + 280x + 140^2 - 340^2 = 0
---
2x^2 + 280x - 96000 = 0
---
x^2 + 140x - 48000 = 0
(x-160)(x+300) = 0
---
Positive solution:
x = 160 miles (eastbound ship)
x+140 = 300 miles (southbound ship)
------------------
Cheers,
Stan H.
==================

How far did the eastbound ship travel? ________ miles
How far did the southbound ship travel? ________ miles